# Effects of Geometric Modelling and Blood Rheology in Patient-Specific Arterial Blood Flow Simulations with Speed-Accuracy Trade-Off Analysis

### v3

@@ 4 Discussion
The study provides two broad observations. First, geometric simplifications can be useful for generating fast but approximate insights, especially when time constraints exist in clinical decision-making. The approximate solutions can often be improved using suitable scale analysis. However, for a detailed analysis, retaining complete geometrical fidelity is essential. Secondly, for larger arteries, the Newtonian model can provide qualitatively similar predictions faster than the non-Newtonian ones in terms of CPU time. The findings also emphasise the importance of combining accurate hemodynamic indicators such as time-averaged wall shear stress and the $Q$-function with newly developed blood viscosity models for better prediction of blood-tissue interaction. These insights could aid in designing personalised treatments, evaluating surgical interventions, and improving diagnostic accuracy in detecting the onset of cardiovascular diseases.
5 Conclusions
@@
A detailed study on a patient-specific geometry of the descending aorta with two simplified representations and three viscosity models is conducted. The geometry is extracted from CT scan data shared by a hospital and reconstructed using image processing tools. Realistic time-dependent boundary conditions are applied at the inlet and outlet in terms of pulsatile velocity (mean Re = 313 and peak Re = 1687) and outlet pressure, respectively. For the geometrical approximation, the original artery (termed Level 1) is further simplified into two more levels, denoted as Level 2 and Level 3. Level 2 is based on the cross-sections of Level 1, which is divided into 25 sections, with each part having a representative average diameter. Thus, Level 2 captures some of the undulations and area fluctuations of Level 1 but loses the local asymmetry of the original. Level 3 represents a smooth duct based on the inlet and outlet cross-sectional area of the actual artery. The 3D simulations are performed with three blood rheology models – Newtonian, Carreau-Yasuda, and Apostolidis and Beris. Various aspects of blood flow for all the levels of the descending aorta geometry (with selected blood viscosity models) are studied. The following conclusions are arrived at in the present work.
@@
    As the geometrical complexity is averaged out, the overall flow field becomes uniform, with a reduction in the strength of the recirculation zones near wall perturbations. For Level 1 geometry, the flow field shows a significant number of recirculation zones near the walls. Some of these zones are visible in Level 2 as well, though with reduced intensity. No such zone of reversed flow is visible in Level 3, where the flow field is practically unidirectional.
@@
    The simplification of the geometry affects the strain rate distribution, further influencing the wall shear stress (WSS). Simulations show that Level 2 geometry could approximate the WSS distribution reasonably well, capturing major patterns and magnitudes while missing the local asymmetry in the stress profiles. However, Level 3 shows only a smooth variation of the WSS without showing undulations.
@@
    The correlation analysis indicates that Level 2 prediction of WSS maintain a strong agreement with Level 1 results, thereby preserving greater accuracy when employing non-Newtonian models. In contrast, this accuracy significantly diminishes in Levels 2 and 3 when using Newtonian viscosity. Pressure remains well-correlated except during the diastolic phases C and D for the Newtonian model.
@@
    The variations of the hemodynamic indicators such as TAWSS and OSI bolster these earlier assertions further. Level 2 predictions are particularly close to those obtained in Level 1, especially when both sets of data are averaged over the circumference. However, Level 3 predictions are without any fluctuations and are, at best, a representative average.
@@
    The speed test showed that computations with Levels 2 and 3 could be about 4 times faster than Level 1. The Newtonian model needs about 10 times fewer iterations than the Apostolidis and Beris model for a comparable convergence level and grid refinement.
@@ Acknowledgment
Authors acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ’PARAM Sanganak’ at IIT Kanpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. Also, we would like to thank the computer centre (www.iitk.ac.in/cc) at IIT Kanpur for providing the resources to carry out the reported work. The authors gratefully acknowledge the kind support provided by the DST NSM project with DST/NSM/R&D_HPC_Applications/2021/18 sanction number .
@@
The authors have no conflicts to disclose.

### v4

@@ IV Discussion
The study provides two broad observations. First, geometric simplifications can help generate fast but approximate insights, primarily when time constraints exist in clinical decision-making. The approximate solutions can often be improved using suitable scale analysis. However, for a detailed analysis, retaining complete geometrical fidelity is essential. Secondly, for larger arteries, the Newtonian model can provide qualitatively similar predictions faster than the non-Newtonian ones in terms of CPU time. The findings also emphasize the importance of combining accurate hemodynamic indicators, such as time-averaged wall shear stress and the $Q$-function, with newly developed blood viscosity models for better prediction of blood-tissue interaction. These insights could aid in designing personalised treatments, evaluating surgical interventions, and improving diagnostic accuracy in detecting the onset of cardiovascular diseases.
V Conclusions
@@
A detailed study of a patient-specific geometry of the descending aorta, with two simplified representations and three viscosity models, is conducted. The geometry is extracted from CT scan data shared by a hospital and reconstructed using image processing tools. Realistic time-dependent boundary conditions are applied at the inlet and outlet, respectively, in terms of pulsatile velocity (mean Re = 313 and peak Re = 1687) and outlet pressure. For the geometrical approximation, the original artery (termed Level 1) is further simplified into two more levels, denoted as Level 2 and Level 3. Level 2 is based on the cross-sections of Level 1, which is divided into 25 sections, with each part having a representative average diameter. Thus, Level 2 captures some of the undulations and area fluctuations of Level 1 but loses the local asymmetry of the original. Level 3 represents a smooth duct based on the inlet and outlet cross-sectional area of the actual artery. The 3D simulations are performed using three blood rheology models: Newtonian, Carreau-Yasuda, and Apostolidis and Beris. Various aspects of blood flow for all the levels of the descending aorta geometry (with selected blood viscosity models) are studied. The following conclusions are drawn in this work.
@@
    As the geometrical complexity is averaged out, the overall flow field becomes uniform, with a reduction in the strength of the recirculation zones near wall perturbations. For Level 1 geometry, the flow field shows a significant number of recirculation zones near the walls. Some of these zones are also visible in Level 2, albeit with reduced intensity. No such zone of reversed flow is visible in Level 3, where the flow field is practically unidirectional.
@@
    The simplification of the geometry affects the strain rate distribution, which in turn influences the wall shear stress (WSS). Simulations indicate that Level 2 geometry can reasonably approximate the WSS distribution, capturing major patterns and magnitudes while missing the local asymmetry in the stress profiles. However, Level 3 shows only a smooth variation of the WSS without showing undulations.
@@
    The correlation analysis indicates that Level 2 prediction of WSS maintains a strong agreement with Level 1 results, thereby preserving greater accuracy when employing non-Newtonian models. In contrast, this accuracy significantly diminishes in Levels 2 and 3 when using Newtonian viscosity. Pressure remains well-correlated, except during phases C and D of the diastolic cycle, for the Newtonian model.
@@
    The variations in hemodynamic indicators, such as TAWSS and OSI, further bolster these earlier assertions. Level 2 predictions are particularly close to those obtained in Level 1, especially when both sets of data are averaged over the circumference. However, Level 3 predictions are without any fluctuations and are, at best, a representative average.
@@
    The speed test revealed that computations with Levels 2 and 3 were approximately four times faster than those with Level 1. The Newtonian model requires approximately 10 times fewer iterations than the Apostolidis and Beris model to achieve a comparable convergence level and grid refinement.
@@ Acknowledgment
Authors acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ’PARAM Sanganak’ at IIT Kanpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. Additionally, we would like to thank the Computer Centre (www.iitk.ac.in/cc) at IIT Kanpur for providing the resources to carry out the reported work. The authors gratefully acknowledge the kind support provided by the DST NSM project with DST/NSM/R&D_HPC_Applications/2021/18 sanction number .
@@
The authors declare that they have no conflicts of interest.